# Optimal Bias-Correction and Valid Inference in High-Dimensional Ridge Regression: A Closed-Form Solution

### v1

@@
Zhaoxing Gao Center for Data Science, Zhejiang University. Email: zhaoxing_gao@zju.edu.cn. I gratefully acknowledge the research support from the NSFC. I thank Ruey S. Tsay for thoughtful and friendly guidance throughout the project, and for encouraging me to develop these results into a paper.
(May 1, 2024)
Ridge regression is an indispensable tool in big data econometrics but suffers from bias issues affecting both statistical efficiency and scalability. We introduce an iterative strategy to correct the bias effectively when the dimension $p$ is less than the sample size $n$. For $p>n$, our method optimally reduces the bias to a level unachievable through linear transformations of the response. We employ a Ridge-Screening (RS) method to handle the remaining bias when $p>n$, creating a reduced model suitable for bias-correction. Under certain conditions, the selected model nests the true one, making RS a novel variable selection approach. We establish the asymptotic properties and valid inferences of our de-biased ridge estimators for both $p<n$ and $p>n$, where $p$ and $n$ may grow towards infinity, along with the number of iterations. Our method is validated using simulated and real-world data examples, providing a closed-form solution to bias challenges in ridge regression inferences.
Keywords: Ridge Regression, Bias Correction, High-Dimension, Ridge Screening, Inference, Bias-Variance Trade-off
@@ 1   Introduction
Ridge regression, or more formally $\ell_{2}$-regularization estimation, is a fundamental tool in econometrics, statistics, and machine learning with applications in various fields of science, technology, engineering, mathematics, medicine, social sciences, and humanities. The idea of $\ell_{2}$-regularization appeared in the early 1940s for the stability of inverse problems; see Tikhonov (1943). It was first introduced to data analysis by Hoerl (1959) and later formulated in Hoerl and Kennard (1970b, a) for providing a robust solution to some of the persistent challenges encountered in traditional linear regression techniques; see Hoerl (1985) for a nice review. Emerging as a fundamental technique in predictive modeling, ridge regression addresses issues such as multicollinearity and overfitting, which commonly afflict predictive models dealing with high-dimensional data. Since its inception, ridge regression’s practical adoption persists due to its superior performance over the least-squares estimator in various scenarios, evident in applications across neuroscience, chemistry, biology, and economics; see Leonard et al. (2023), Zahrt et al. (2019), Otwinowski and Plotkin (2014), Giannone et al. (2021), and Abadie and Kasy (2019), among others, underscoring its empirical effectiveness. From a shrinkage perspective, the ridge estimator also dominates the least-squares solutions in the sense that its mean-squared errors (MSEs) can be smaller, which provides a reasonable explanation on the empirical effectiveness of ridge estimators. See Theobald (1974), Athey and Imbens (2019), Hastie (2020), Hansen (2022a), and a comprehensive introduction to ridge regression in van Wieringen (2023).
@@
The ridge estimator offers a closed-form expression that simplifies both theoretical and empirical analyses. It aligns with the dense modeling techniques of Giannone et al. (2021), which acknowledge the potential significance of all explanatory variables for prediction. Empirical studies, such as those in Giannone et al. (2021), indicate that dense models generally tend to outperform the sparse ones in out-of-sample economic prediction performance. Similarly, Abadie and Kasy (2019) find that the ridge estimators dominate the lasso and the pre-testing estimators in terms of the risks when the effects of different predictors on the dependent variable are “smoothly distributed”. These results suggest that ridge estimators indeed constitute a crucial tool in econometric modeling and economic forecasting, especially in the big data era.
@@
However, as highlighted in Section 2.8 of Athey and Imbens (2019), constructing valid confidence intervals remains a challenge for many regularized methods, including ridge regression, even in asymptotic settings. This long-standing challenge in ridge-type regression involves at least two critical aspects within a linear regression framework: (a) performing hypothesis tests on specific linear combinations of the regression coefficients using the ridge estimators; and (b) deriving confidence or prediction intervals based on the ridge estimators in empirical applications. The primary reason for these challenges is that the inherent bias of ridge estimators poses significant challenges, compromising both statistical efficiency and scalability across various applications. To date, the feasibility of conducting statistical inferences and hypothesis testing on ridge-type estimators without imposing additional structural constraints remains largely unexplored in the literature. This complexity arises from the ridge estimator’s intrinsic bias, which complicates direct statistical inferences despite its elegant closed-form expression. As a result, research addressing the inference challenges of ridge regression in high-dimensional settings is limited, leading to its widespread application across disciplines without comprehensive theoretical investigations.
@@
To the best of our knowledge, there are only a few works in the literature concerning the bias and inference of ridge estimators under different scenarios. Under a sparse structure, Shao and Deng (2012) proposed a threshold ridge regression method and proved its consistency. The method therein actually estimates the projected coefficient vector rather than the true one in the linear model. Dobriban and Wager (2018) derived the limit of the predictive risk of ridge regression and regularized discriminant analysis in a dense random effects model. Bühlmann (2013) proposed to use the lasso to correct the bias of ridge estimators. However, the estimation depends on the existence of an initial estimator which needs to be accurate enough. Zhang and Politis (2022) adopted a similar approach as that in Shao and Deng (2012) and proposed a threshold ridge regression and a bootstrap method to make inferences. However, the method therein still estimates the projected coefficient vector rather than estimating the true one, and the remaining bias may not be asymptotically negligible in general.
@@
To further combat the remaining bias in the de-biased ridge estimators when $p>n$, we introduce a novel ridge-screening (RS) method for covariate selection prior to applying our bias correction procedure. The RS approach constructs a restricted model that inherently encompasses the true model as a subset. This is based on the assumption that only a subset of the covariates holds significance in the linear model. Specifically, we postulate that the number of significant covariates should be less than the sample size, which can appropriately diverge relative to the dimension $p$ and sample size $n$. Crucially, we demonstrate that under certain mild conditions, the true model is inherently nested within the selected one, establishing RS as a novel variable selection approach that offers independent value beyond bias correction. Leveraging this restricted model, our bias-correction procedure can be further applied to the restricted model and effectively rectifies the bias in the resulting ridge estimators
@@
The derivation of the methodology and theory is mainly based on a fixed design, which is similar to the setting in Hansen (2022b), and the results throughout this paper are valid for random regressors by conditioning on the design matrix. This paper rigorously establishes the asymptotic properties and provides valid inferences of our estimators for both $p<n$ and $p>n$ under some relaxed and intuitive conditions. Furthermore, we delve into the bias-variance trade-off of our de-biased ridge estimators, examining its relationship with the number of iterations in the bias-correction procedure. To validate our methodology, we provide empirical evidence using both simulated and real-world data. The prediction intervals constructed by the proposed method are indeed satisfactory for forecasting the U.S. macroeconomic series using factor-augmented regression. Moreover, the RS method can further enhance the coverage rates of these prediction intervals. These empirical validations highlight the practical efficacy and adaptability of our approach across a wide range of high-dimensional regression settings.
@@
The contributions of this paper are multi-fold. First, the proposed bias-correction method is simple and easy to implement. In fact, the proposed approach is a systematic procedure and the resulting de-biased estimator has a closed-form expression. Second, the optimality of the proposed bias-correction procedure consists of two aspects: (a) it achieves complete bias correction when the covariate dimension $p$ is smaller than the sample size $n$; and (b) the remaining bias of the de-biased ridge estimator in the scenario when $p>n$ is unattainable through any linear transformations of the response vector. These results distinguish our work from the existing ones that only part of the bias can be corrected in most of the aforementioned literature. Third, we also propose a novel variable selection procedure, namely the ridge-screening (RS) method, which screens out some insignificant variables based on the de-biased ridge estimator due to its optimality. Fourth, we also establish the asymptotic properties of the de-biased ridge estimators in both scenarios when $p<n$ and $p>n$, and it is shown that the de-biased ridge estimators are asymptotically normal, which provides valid inference methods for the ridge estimators. Fifth, we develop a procedure to construct confidence and prediction intervals in ridge regressions using our proposed de-biased estimators and associated inference methods. Finally, we establish the bias-variance trade-off of our proposed approach both theoretically and through validation with simulated data. It’s important to note that, unlike the scenario described in Section 2.8 of Athey and Imbens (2019) where many inference approaches for regularized machine learning methods compromise predictive performance, our proposed procedure focuses on correcting the bias and rendering the estimators suitable for inferences, without adversely affecting their predictive performance.
@@
We highlight that the asymptotic framework adopted in this paper is slightly different from traditional approaches in the literature. Typically, in conventional frameworks, the asymptotic properties are established when the dimension and/or the sample size are approaching infinity. However, in our study, most of the asymptotic results are derived under a different scenario that the number of iterations in the bias-correction procedure tends towards infinity for any given configuration of the dimension $p$ and the sample size $n$. This configuration can be sufficiently large to encompass the framework of big data analysis. One of the primary motivations behind this choice is to demonstrate the validity of our proposed procedure by explicitly providing the exact bias and covariance terms of the de-biased estimators in this paper. This approach is also reasonable because it reflects common scenarios encountered in practical data analysis, where datasets often have fixed dimensions and sample sizes. Without this setting, the dimensions of the bias and covariance terms of the de-biased estimators would expand to infinity if we considered $n$ and $p$ approaching infinity, making them challenging to formulate and describe theoretically. Moreover, our asymptotic results remain valid as $n$ and $p$ approach infinity. This holds true so long as we initially allow the number of iterations to increase towards infinity at a moderate rate. In this asymptotic manner, we can symbolically retain the forms of the bias and covariance terms for a growing configuration of $(p,n)$. Therefore, it’s important to emphasize that our chosen asymptotic framework does not undermine the validity of the proposed bias-correction procedures.
@@
The rest of the paper is organized as follows: Section 2 introduces the ridge estimation and its bias-correction procedure in scenarios when $p<n$ and $p>n$, where a ridge-screening method is also introduced for variable selection. Following this, Section 2 also presents the inference methodologies for the proposed de-biased estimators. Section 3 examines the proposed approach’s finite-sample performance via Monte-Carlo simulations. Section 4 provides an empirical application of the proposed method, and Section 5 offers conclusive insights. All proofs and derivations for the asymptotic results are available in an online Appendix.
@@
Notation: We use the following notation. For a $p\times 1$ vector ${\mathbf{u}}=(u_{1},...,u_{p})^{\prime}$, $\|{\mathbf{u}}\|_{2}=\sqrt{\sum_{i=1}^{p}|u_{i}|^{2}}$ is the $\ell_{1}$-norm and $\|{\mathbf{u}}\|_{\infty}=\max_{1\leq i\leq p}|u_{i}|$ is the $\ell_{\infty}$-norm. ${\mathbf{I}}_{p}$ denotes the $p\times p$ identity matrix. For a matrix ${\mathbf{H}}$, its operator norm is $\|{\mathbf{H}}\|_{2}=\sqrt{\lambda_{\max}({\mathbf{H}}^{\prime}{\mathbf{H}})}$, where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue of a matrix, and $\|{\mathbf{H}}\|_{\min}$ is the square root of the minimum non-zero eigenvalue of ${\mathbf{H}}{\mathbf{H}}^{\prime}$. $|{\mathbf{H}}|$ denotes the absolute value of ${\mathbf{H}}$ elementwisely. The superscript denotes the transpose of a vector or matrix. We also use the notation $a\asymp b$ to denote $a=O(b)$ and $b=O(a)$ or $a$ and $b$ have the same order.
@@ 2.1 High-dimensional Linear Regression
Let $\{({\mathbf{x}}_{1},y_{1}),...,({\mathbf{x}}_{n},y_{n})\}$ be a given sample of centered observable data. We consider the problem of estimating a $p$-dimensional vector $\boldsymbol{\beta}$ from the following linear model:
@@
Note that the least-squares estimator $\widehat{\boldsymbol{\beta}}_{lse}$ is only well-defined if $({\mathbf{X}}^{\prime}{\mathbf{X}})^{-1}$ exists. In a high-dimensional setting, if the columns of the design matrix ${\mathbf{X}}$ are linearly dependent, for example, this is obviously true when $p>n$, this collinearity among the columns implies that ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is singular, rendering $\widehat{\boldsymbol{\beta}}_{lse}$ an invalid estimator. To make the least-squares estimator in (3) a well-defined quantity, we modify the definition in (3) as
@@
where $({\mathbf{X}}^{\prime}{\mathbf{X}})^{+}$ is the Moore-Penrose generalized inverse. It is not hard to see that the estimator in (4) reduces to the one in (3) if $p<n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible. Therefore, we will denote the estimator in (4) as the least-squares solution throughout this article.

### v2

@@
Zhaoxing Gao1 and Ruey S. Tsay2111Corresponding author: ruey.tsay@chicagobooth.edu (R.S. Tsay). Booth School of Business, University of Chicago, 5807 S. Woodlawn Avenue, Chicago, IL 60637, USA.
1University of Electronic Science and Technology of China
2University of Chicago
Ridge regression is an indispensable tool in big data analysis. Yet its inherent bias poses a significant and longstanding challenge, compromising both statistical efficiency and scalability across various applications. To tackle this critical issue, we introduce an iterative strategy to correct bias effectively when the dimension $p$ is less than the sample size $n$. For $p>n$, our method optimally mitigates the bias such that any remaining bias in the proposed de-biased estimator is unattainable through linear transformations of the response data. To address the remaining bias when $p>n$, we employ a Ridge-Screening (RS) method, producing a reduced model suitable for bias correction. Crucially, under certain conditions, the true model is nested within our selected one, highlighting RS as a novel variable selection approach. Through rigorous analysis, we establish the asymptotic properties and valid inferences of our de-biased ridge estimators for both $p<n$ and $p>n$, where, both $p$ and $n$ may increase towards infinity, along with the number of iterations. We further validate these results using simulated and real-world data examples. Our method offers a transformative solution to the bias challenge in ridge regression inferences across various disciplines.
Keywords: Bias Correction, High-Dimension, Inference, Ridge Regression, Ridge Screening
@@ 1   Introduction
Ridge regression, or more formally $\ell_{2}$-regularisation estimation, is a fundamental tool in econometrics, statistics, and machine learning with applications in many fields of science, technology, engineering, mathematics, medicine, social sciences, and humanities. The idea of $\ell_{2}$-regularisation appeared in the early 1940s for the stability of inverse problems; see Tikhonov (1943). It was first introduced to data analysis by Hoerl (1959) and later formulated in Hoerl and Kennard (1970b, a) for providing a robust solution to some of the persistent challenges encountered in traditional linear regression techniques; see Hoerl (1985) for a nice review. Emerging as a fundamental technique in predictive modelling, ridge regression addresses issues such as multicollinearity and overfitting, which commonly afflict predictive models dealing with high-dimensional data. Since its inception, ridge regression’s practical adoption persists due to its superior performance over the least-squares estimator in various scenarios, evident in applications across neuroscience, chemistry, biology, and economics; see Leonard et al. (2023), Zahrt et al. (2019), Otwinowski and Plotkin (2014), Giannone et al. (2021), and Abadie and Kasy (2019), among others, underscoring its empirical effectiveness. From a shrinkage perspective, the ridge estimator also dominates the least-squares solutions in the sense that its mean-squared errors (MSEs) can be smaller, which provides a reasonable explanation on the empirical effectiveness of ridge estimators. See Theobald (1974), Athey and Imbens (2019), Hastie (2020), Hansen (2022a), and a comprehensive introduction of ridge regression in van Wieringen (2023).
@@
The ridge estimator offers a closed-form expression that simplifies both theoretical and empirical analyses. It aligns with the dense modelling techniques of Giannone et al. (2021), which acknowledge the potential significance of all explanatory variables for prediction. Empirical studies, such as those in Giannone et al. (2021), indicate that dense models generally tend to outperform the sparse ones in out-of-sample economic prediction. Similarly, Abadie and Kasy (2019) find that the ridge estimators dominate the Lasso and the pre-testing estimators in terms of risks when the effects of different predictors on the dependent variable are “smoothly distributed”. These results suggest that ridge estimators indeed constitute a crucial tool in statistical modelling and economic forecasting, especially in the big data era.
@@
However, as highlighted in Section 2.8 of Athey and Imbens (2019), constructing valid confidence intervals remains a challenge for many regularised methods, including ridge regression, even in asymptotic settings. This long-standing challenge in ridge-type regression involves at least two critical aspects within a linear regression framework: (a) performing hypothesis testing on specific linear combinations of the regression coefficients using the ridge estimators; and (b) deriving confidence or prediction intervals based on the ridge estimators in empirical applications. The primary reason for these challenges is that the inherent bias of ridge estimators poses significant challenges, compromising both statistical efficiency and scalability across various applications. To date, the feasibility of conducting statistical inferences and hypothesis testing on ridge-type estimators without imposing additional structural constraints remains largely unexplored in the literature. This complexity arises from the ridge estimator’s intrinsic bias, which complicates direct statistical inferences despite its elegant closed-form expression. As a result, research addressing the inference challenges of ridge regression in high-dimensional settings is limited, leading to its widespread application across disciplines without comprehensive theoretical investigations.
@@
To the best of our knowledge, there are only a few works in the literature concerning the bias and inference of ridge estimators under different scenarios. Assuming a sparse structure, Shao and Deng (2012) proposed a threshold ridge regression method and proved its consistency. The method therein actually estimates the projected coefficient vector rather than the true one in the linear model. Dobriban and Wager (2018) derived the limit of the predictive risk of ridge regression and regularised discriminant analysis in a dense random effects model. Bühlmann (2013) proposed to use the Lasso to correct the bias of ridge estimators. However, the estimation depends on the existence of an initial estimator which must be sufficiently accurate. Zhang and Politis (2022) adopted a similar approach as that in Shao and Deng (2012) and proposed a threshold ridge regression and a bootstrap method to make inferences. However, the method therein still estimates the projected coefficient vector rather than the true one, and the remaining bias may not be asymptotically negligible in general.
@@
To further combat the remaining bias in the de-biased ridge estimators when $p>n$, we introduce a novel ridge-screening (RS) method for selecting covariates prior to applying our bias correction procedure. The RS approach constructs a restricted model that inherently encompasses the true model as a subset. This is based on the assumption that only a subset of the covariates holds significance in the linear model. Specifically, we postulate that the number of significant covariates should be less than the sample size, which can appropriately diverge relative to the dimension $p$ and sample size $n$. Crucially, we demonstrate that under certain mild conditions, the true model is inherently nested within the selected one, establishing RS as a novel variable selection approach that offers additional value beyond bias correction. Leveraging this restricted model, our bias-correction procedure can be further applied to the restricted model and effectively rectifies the bias in the resulting ridge estimators
@@
The derivation of the proposed methodology and theory is mainly based on a fixed design, which is similar to the setting in Hansen (2022b), but the results throughout this paper remain valid for random regressors by conditioning on the design matrix. Thus, this paper rigorously establishes asymptotic properties and provides valid inferences of the proposed bias-corrected estimators for both $p<n$ and $p>n$ under some relaxed and intuitive conditions. Furthermore, we delve into the bias-variance trade-off of our de-biased ridge estimators, examining its relationship with the number of iterations in the bias-correction procedure. To validate our methodology, we provide empirical evidence using both simulated and real-world data. The prediction intervals constructed by the proposed method are indeed satisfactory for forecasting the U.S. macroeconomic series using factor-augmented regression. Moreover, the RS method can further enhance the coverage rates of these prediction intervals. These empirical validations highlight the practical efficacy and adaptability of the proposed approach across a wide range of high-dimensional regression settings.
@@
The contributions of this paper are multi-fold. First, the proposed bias-correction method is simple and easy to implement. In fact, the proposed approach is a systematic procedure and the resulting de-biased estimator has a closed-form expression. Second, the optimality of the proposed bias-correction procedure consists of two aspects: (a) it achieves complete bias correction when the covariate dimension $p$ is smaller than the sample size $n$; and (b) the remaining bias of the de-biased ridge estimator in the scenario when $p>n$ is unattainable through any linear transformations of the response vector. These results distinguish our work from the existing ones that only part of the bias can be corrected in most of the aforementioned literature. Third, we also propose a novel variable selection procedure, namely the ridge-screening (RS) method, which screens out some insignificant variables based on the de-biased ridge estimator due to its optimality. Fourth, we establish certain asymptotic properties of the de-biased ridge estimators in both scenarios when $p<n$ and $p>n$, and it is shown that the de-biased ridge estimators are asymptotically normal, which provides valid inference methods for the ridge estimators. Fifth, we develop a procedure to construct confidence and prediction intervals in ridge regressions using the proposed de-biased estimators and associated inference methods. Finally, we establish the bias-variance trade-off of the proposed approach both theoretically and through validation with simulated data. It’s important to note that, unlike the scenario described in Section 2.8 of Athey and Imbens (2019) where many inference approaches for regularised machine learning methods compromise predictive performance, our proposed procedure focuses on correcting the bias and rendering the estimators suitable for statistical inferences, without adversely affecting their predictive performance.
@@
We highlight that the asymptotic framework adopted in this paper is slightly different from traditional approaches in the literature. Typically, in conventional frameworks, the asymptotic properties are established when the dimension and/or the sample size are approaching infinity. However, in our study, most of the asymptotic results are derived under a different scenario that the number of iterations in the bias-correction procedure tends to infinity for any given configuration of the dimension $p$ and the sample size $n$. This configuration can be sufficiently large to encompass the framework of big data analysis. One of the primary motivations behind this choice is to demonstrate the validity of the proposed procedure by explicitly providing the exact bias and covariance terms of the de-biased estimators in this paper. This approach is also reasonable because it reflects common scenarios encountered in empirical data analysis, where datasets often have fixed dimensions and sample sizes. Without this setting, the dimensions of the bias and covariance terms of the de-biased estimators would expand to infinity if we considered $n$ and $p$ approaching infinity, making them challenging to formulate and describe theoretically. Moreover, our asymptotic results remain valid as $n$ and $p$ approach infinity. This holds true so long as we initially allow the number of iterations to increase towards infinity at a moderate rate. In this asymptotic manner, we can symbolically retain the forms of the bias and covariance terms for a growing configuration of $(p,n)$. Therefore, it’s important to emphasise that our chosen asymptotic framework does not undermine the validity of the proposed bias-correction procedures.
@@
The rest of the paper is organised as follows: Section 2 introduces the ridge estimation and its bias-correction procedure in scenarios when $p<n$ and $p>n$, where a ridge-screening method is also introduced for variable selection. Section 2 presents the inference methodologies for the proposed de-biased estimators. Section 3 examines the finite-sample performance pf the proposed method via Monte-Carlo simulations. Section 4 provides an empirical application of the proposed method, and Section 5 offers some concluding insights. All proofs and derivations for the asymptotic results are available in an online Supplementary Material.
@@
Notation: We use the following notation. For a $p\times 1$ vector ${\mathbf{u}}=(u_{1},...,u_{p})^{\prime}$, $\|{\mathbf{u}}\|_{2}=\sqrt{\sum_{i=1}^{p}|u_{i}|^{2}}$ is the $\ell_{1}$-norm and $\|{\mathbf{u}}\|_{\infty}=\max_{1\leq i\leq p}|u_{i}|$ is the $\ell_{\infty}$-norm. ${\mathbf{I}}_{p}$ denotes the $p\times p$ identity matrix. For a matrix ${\mathbf{H}}$, its operator norm is $\|{\mathbf{H}}\|_{2}=\sqrt{\lambda_{\max}({\mathbf{H}}^{\prime}{\mathbf{H}})}$, where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue of a matrix, and $\|{\mathbf{H}}\|_{\min}$ is the square root of the minimum non-zero eigenvalue of ${\mathbf{H}}{\mathbf{H}}^{\prime}$. $|{\mathbf{H}}|$ denotes the absolute value of ${\mathbf{H}}$ elementwisely. The superscript denotes the transpose of a vector or matrix. We also use the notation $a\asymp b$ to denote $a=O(b)$ and $b=O(a)$ or $a$ and $b$ are of the same order.
@@ 2.1 High-dimensional Linear Regression
Let $\{({\mathbf{x}}_{1},y_{1}),...,({\mathbf{x}}_{n},y_{n})\}$ be a given sample of centred observable data. We consider the problem of estimating a $p$-dimensional vector $\boldsymbol{\beta}$ from the following linear model:
@@
This least-squares estimator $\widehat{\boldsymbol{\beta}}_{lse}$ is only well-defined if $({\mathbf{X}}^{\prime}{\mathbf{X}})^{-1}$ exists. In a high-dimensional setting, if the columns of the design matrix ${\mathbf{X}}$ are exactly linearly dependent, for example, this is obviously true when $p>n$, this collinearity among the columns implies that ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is singular, rendering $\widehat{\boldsymbol{\beta}}_{lse}$ an invalid estimator. To make the least-squares estimator in (3) a well-defined quantity, we modify the definition in (3) as
@@
where $({\mathbf{X}}^{\prime}{\mathbf{X}})^{+}$ is the Moore-Penrose generalised inverse. It is not hard to see that the estimator in (4) reduces to the one in (3) if $p<n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible. Therefore, we will denote the estimator in (4) as the least-squares solution throughout this article.